OpenStudy (exoexoexo):

deshawn and luke had an equal number of stickers. after deshawn used 26 stickers and luke lost 38 stickers, deshawn had twice as many stickers as luke. how many stickers did each of them have in the beginning?

OpenStudy (anonymous):

You can write this as a system of equations. Deshawn started with x stickers: x-26=2y Luke started with x stickers: x-38=y Multiply the second equation by 2, getting 2x-76=2y Then, you can write 2x-76=x-26 Using the Addition Property of Equality, you can add 76 to each side making it 2x=x+50 Then, use the Subtraction Property and get rid of one x, making the final equation x=50 Therefore, each boy had 50 stickers to start with Hope this helps!